OpenStudy (anonymous):

Can someone check my answer i got C.) please help me Solve for x in the following equation 5x^2-56=0 A.) x= -7 ; 8/5 B.)x=-7 ;8/5 C.) x=7 ;-8/5 D.)x=7;8/5

OpenStudy (anonymous):

You sure you typed the problem statement correctly? That should be x = ±sqrt(56/5)

OpenStudy (anonymous):

no typed it rigth

OpenStudy (anonymous):

Well, none of the answers are correct. You don't even need any fancy quadratic methods for the problem as it's given. 5x^2 - 56 = 0 5x^2 = 56 x^2 = 56/5 x = ±sqrt(56/5)

OpenStudy (anonymous):

...so roughly ±3.35